I don't understand this....
Find the numbers of each of the following polynomials for the given values of the variable.
-x2 – 5x – 6, x = -3 and x = -2
-x(2 is and exponent)sorry not trying to confuse you.

Answer by adamchapman(301)

When x=-3,
-x^2–5x–6=-9+15-6=0
When x=-2,
-x^2–5x–6=-4+10-6=0
I hope this helps
P.S. I am trying to start up my own homework help website. I would be extremely grateful if you would e-mail me some feedback on the help you received to adam.chapman@student.manchester.ac.uk
Adam